Sasha owns two investments, A and B, that have a combined total value of 42,000 dollars. Investment A is expected to pay 28,200 dollars in 1 year(s) from today and has an expected return of 12.25 percent per year. Investment B is expected to pay 21,431 in 2 years from today and has an expected return of R per year. What is R, the expected annual return for investment B? Answer as a rate in decimal format so that 12.34% would be entered as .1234 and 0.98% would be entered as .0098.
Answer #1

We need to equate the present value of both the investments to the combined total value.

Present value =Future value/(1+Rate)^number of periods

Hence

42000 = 28200/(1+0.1225)^1 + 21431/ (1+R)^2

42000 = 25122.49+ 21431/ (1+R)^2

21431/ (1+R)^2 = 16877.51

(1+R)^2 = 1.269796648

1+R = (1.269796648)^(1/2)

1+R =1.126852541

R = 0.1269